OpenStudy (anonymous):

the trust of a propeller varies jonly as the square of the number n of revolutions per secondand the fourth power of its diameter d. a) set up the proportional relationship b) what is the effect on T if n is doubled and d is halved?

OpenStudy (anonymous):

Thrust/(n^2 *d^4) T/(n^2 *d^4) Choose values for n and d. Let n=3 and d=8 T/( 3^2 * 8^4)= T/(9*4,096)= T/36,864 Let T=36,864 Therefore we use this value for T and see what happens. Now use the conditions: n is doubled and d is halved. let n = 3*2= 6 and d= 8/2 =4. Plug in new values: T/(6^2 * 4^4)= T/(36 * 256)= T/9216 36,864/9216= 4 Therefore the Thrust is increased 4 times.

OpenStudy (anonymous):

Mertsj are you there